OpenStudy (across):

Handle with care:\[\frac{\partial u}{\partial t}+c\frac{\partial u}{\partial x}+au=0.\]Let us solve for u(t,x)!

OpenStudy (across):

I actually want to re-do this to refresh my memory.

OpenStudy (across):

If I recall correctly, we need to perform a substitution. That is,\[\xi=x-ct.\]

OpenStudy (across):

This is so that the equation can be simplified by having the plane be moving along with the "wave" being described. Something akin to Einstein's relativistic concepts.

OpenStudy (across):

\[u(t,x)=v(t,x-ct)=v(t,\xi).\]We now need to find how to express the partial differentials in terms of the new function v.

OpenStudy (anonymous):

Would separation of variables work? \[(af(x)+cf'(x))/f(x) = - g'(t)/g(t), u(x, t) = g(t)*f(x)\]

OpenStudy (anonymous):

Or only for the heat equation?

OpenStudy (across):

\[\frac{\partial u}{\partial t}=\frac{\partial v}{\partial t}-c\frac{\partial v}{\partial \xi}\]\[\frac{\partial u}{\partial x}=\frac{\partial v}{\partial \xi}\]

OpenStudy (across):

Hmm, let me think about it.

OpenStudy (across):

I'm a bit confused with what you're trying to say with that equation up there. Is that, on the very right, the convolution of those two functions?

OpenStudy (anonymous):

No, multiplication.

OpenStudy (anonymous):

Substituting you can get \[\frac{\partial v}{\partial t} = -au\]

OpenStudy (across):

Ah, I see what you're doing now!

OpenStudy (across):

We could do\[\frac{\partial v}{\partial t}+av=0\]\[\frac{\partial}{\partial t}\left [ e^{at}v \right ]=0\]Since it's now basically an ODE.

OpenStudy (across):

\[e^{at}v=f(\xi)\]\[v(t,\xi)=f(\xi)e^{-at}\]

OpenStudy (anonymous):

I'm not sure about that because the RHS has the original function \[u\] and the LHS has modified version \[v\] with the coordinate transformation.

OpenStudy (across):

Can't we replace the u on the RHS with the new function v?

OpenStudy (across):

If we can, then by re-substituting we reach\[u(t,x)=f(x-ct)e^{-at}\]to be a solution, which I'm not too sure about. I'll recheck.

OpenStudy (anonymous):

I substituted that back into the original equation and it checked out for me

OpenStudy (across):

Phew! I got carried away reviewing all my notes just now; it's quite an intriguing subject!

OpenStudy (anonymous):

\[\frac{\partial u}{\partial t} = -(af(x - ct)e^{-at} + cf'(x - ct)e^{-at})\] and \[\frac{\partial u}{\partial x} = f'(x - ct)e^{-at}\]

OpenStudy (anonymous):

It really is! That's why I want to go to school for Math/CS

OpenStudy (across):

Wait, are you not in college? o.o

OpenStudy (anonymous):

Just a lowly high school student. But I enjoy math.

OpenStudy (across):

Oh my! I am a graduate student. Haha, you must be really, really smart!

OpenStudy (anonymous):

Lol, Well thank you. I'm sure you are too, I mean graduate school is pretty tough.

OpenStudy (across):

They don't teach anything beyond Calculus I in American high schools (excluding outlier schools). Where are you from?

OpenStudy (anonymous):

I'm from Indiana. I actually taught myself most of what I know about advanced mathematics. Mainly from lecture notes that professors posted from their classes. Right now I'm taking Calculus 1 in school.

OpenStudy (anonymous):

I did take a few college dual credit trigonometry and college algebra classes earlier in high school.

OpenStudy (across):

I see! Long live the open course-ware, huh? :) I wish I would've taught myself more interesting things whilst still a high school student. You're doing a great thing!

OpenStudy (anonymous):

Well, I've always enjoyed learning new things.

OpenStudy (anonymous):

Mathematica for the folks who cannot integrate x with pencil and paper.